OpenStudy (anonymous):

I'm sorry I know this is math but no one's helping me in that forum so please help if you can Simplify 1. 5-3x+y+6 2. 2x+3(x-5)+9 3. -20-(-5)*(-5)[2squared] the bracketed part "2squared" is supposed to be exactly what is says. But, there's no way to put an actual squared number.

OpenStudy (anonymous):

For the first one all you have to do is 5+6. Second one distribute 3 over (x-5) meaning multiply 3 by both x and -5, and collect like terms. When you have exponents you can use this "hat" 2^2 that implies squared, you can also write 2^100 etc, but for that one just used PEDMAS/ BEDMAS what ever suits you. Show us an attempt and we'll go from there, good luck.

OpenStudy (anonymous):

thank you, batman. so first is 11-3x+y and second 5x-15+9? or would it be 5x-24? Is it correct to combine something inside the parenthesis to outside of it? third... Do I really just use PEMDAS?

OpenStudy (anonymous):

First one right Second one 5x-15+9 =>5x-6 that's good Yes ;)

OpenStudy (anonymous):

Thank you so much! But... How do I use pemdas if everything is in parenthesis? It's an odd problem to work with, I don't know where to start on it.

OpenStudy (here_to_help15):

just work with everything thats in the parenthesis as PEMDAS States

OpenStudy (anonymous):

I honestly have no idea what to do for it...

OpenStudy (anonymous):

\[-20-(-5)(-5)*(2^2)\] so first 2^2 = 4 (-5)(-5) = 25, so (25)(4) = 100 -20-100 = ?

OpenStudy (here_to_help15):

Then you mulyiply

OpenStudy (anonymous):

Since they are both in brackets go in order with pedmas, you see exponents is right after parenthesis so go with that since it's the next operation :)

OpenStudy (here_to_help15):

Ok @Sarcasmus i'll help you here if you help me in french?

OpenStudy (anonymous):

Does that make sense?

OpenStudy (here_to_help15):

Parenthesis Exponent Multiplication Division Addition Subtraction you break down a problem in this order.